# Can a Crow Hatch a Falcon? Lineage Matters in Predicting Large Language Model Performance

### v1

@@ Abstract
Accurately forecasting the performance of Large Language Models (LLMs) before extensive fine-tuning or merging can substantially reduce both computational expense and development time. Although prior approaches like scaling laws account for global factors such as parameter size or training tokens, they often overlook explicit lineage relationships—i.e., which models are derived or merged from which parents. In this work, we propose a novel Lineage-Regularized Matrix Factorization (LRMF) framework that encodes ancestral ties among LLMs via a graph Laplacian regularizer. By leveraging multi-hop parent–child connections, LRMF consistently outperforms conventional matrix factorization and collaborative filtering methods in both instance-level and benchmark-level performance prediction. Our large-scale study includes 2,934 publicly available Hugging Face models and 21,000+ instances across 6 major benchmarks, showing that lineage constraints yield up to 7–10 percentage points higher correlation with actual performance compared to baselines. Moreover, LRMF effectively addresses the cold-start problem, providing accurate estimates for newly derived or merged models even with minimal data. This lineage-guided strategy thus offers a resource-efficient way to inform hyperparameter tuning, data selection, and model combination in modern LLM development.
@@ 1 Introduction
If one could accurately predict the performance of a proposed LLM variant—even before running large-scale optimization—it would profoundly reduce the search space. Such foresight would streamline choosing (1) the right model size and data volume for pre-training, (2) fine-tuning hyperparameters and data composition (e.g., domain mix), and (3) whether combining certain model checkpoints is worthwhile. Indeed, prior work has sought to provide partial solutions. Classic scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022) predict an LLM’s loss on the training corpus as a function of parameter size and token count. These laws help developers extrapolate the compute needed to reach a desired perplexity or training loss. Subsequent studies extend these ideas to specific benchmarks, e.g., MMLU (Hendrycks et al., 2021a; Wang et al., 2024), BBH (Suzgun et al., 2023), or MATH (Hendrycks et al., 2021b), by fitting regression models or matrix factorization frameworks to empirical performance data (Zhang et al., 2024; Polo et al., 2024; Kipnis et al., 2025).
@@
We conduct a large-scale empirical study encompassing 2,934 publicly available Hugging Face models and 21,000+ instances across 6 major benchmarks. We demonstrate that incorporating lineage constraints—i.e., explicitly modeling which models were fine-tuned or merged from which parents—consistently boosts predictive accuracy. In fact, compared to standard matrix factorization baselines, our method improves correlation with true performance by up to 7–10 percentage points on certain tasks, affirming that lineage is a key factor in understanding and forecasting LLM behaviors.
Moreover, by leveraging explicit lineage relationships (parent–child or merges), our method LRMF tackles the cold-start problem, enabling accurate performance estimates for newly derived or merged models without costly full re-benchmarking. Indeed, our experiments reveal that even with as few as 10 labeled instances per model, we can achieve near state-of-the-art correlation with fully evaluated baselines—a remarkable improvement over standard collaborative filtering or scaling-based approaches. Taken together, our findings suggest that preserving and exploiting the lineage graph is vital for guiding resource-efficient decisions in LLM development.
@@ 3 Method
Our goal is to predict the performance of Large Language Models (LLMs) on specific tasks or benchmark instances by leveraging partial observations and the inherent relationships among models. We first outline existing approaches based purely on observed performance and factors, then propose our lineage-based method.
@@ 4.1 Experimental Setting
We compared three methods in our experiments:
Baseline: Neural Collaborative Filtering with Factors (NCF with factors).
We used the same network architecture as in previous work (Zhang et al., 2024), consisting of two hidden layers with 128 units each. Due to practical constraints when collecting data from the leaderboard, we limited the model factors to three key attributes: architecture type (50 distinct categories including qwen2forcasualllm, gemma2forcasuallm, etc.), model type (3 categories: Fine-Tuned models, Merged models, and Others including base models), and parameter size. For task factors, we included benchmark identification (6 main benchmarks divided into 39 sub-benchmarks). These factors were encoded as embeddings and concatenated with the model and instance embeddings before being processed by the MLP.
Proposed-1: Model Lineage Averaging.
We constructed lineage relationships by extracting information from Hugging Face model cards. For Fine-Tuned models, we identified parent models from the “base_model:finetune:” tag. For Merged models, we collected parent models from both the “base_model:merge:” tag and the merge_config.json file when available. This approach allowed us to establish direct genealogical connections between derivatives and their parent models, forming the basis for our neighborhood definitions in the lineage averaging process.
Proposed-2: Lineage-Regularized Matrix Factorization.
For our proposed method, we constructed the instance similarity graph using semantic embeddings of the prompts. Specifically, we utilized Snowflake/snowflake-arctic-embed-l-v2.0 to generate high-quality embeddings of the prompt text, and then computed cosine similarities between these embeddings to define edges in the instance graph. For each instance, we retained connections to its top-$k$ most similar instances (with $k=20$ in our experiments). The model similarity graph was constructed using the same lineage information described in the Model Lineage Averaging baseline, where each edge represents a direct derivation relationship between models. The model was implemented in PyTorch and optimized using the Adam optimizer with a learning rate of 0.003. We trained for a maximum of 10,000 epochs with early stopping when no improvement was observed on the validation set for 100 consecutive epochs.
@@
Detailed hyperparameter tuning procedures and their impact on performance are presented in Appendix B. For all methods, we performed grid searches over relevant regularization parameters ($\lambda_{L2}$, $\lambda_{\mathcal{M}}$, $\lambda_{\mathcal{X}}$) and selected the configuration that maximized AUC-ROC on the development set.
@@
For our experiments, we evaluate each method by calculating the correlation coefficient between the estimated performance and the actual performance of LLMs. This metric was chosen because our primary goal is to accurately capture the relative performance ranking among different models rather than precisely matching the absolute performance values. A high correlation coefficient indicates that the method correctly preserves the performance hierarchy among models, which is crucial for identifying promising model variants during development without requiring exhaustive evaluations. We report Pearson correlation coefficients across various benchmarks to assess each method’s predictive capability.
@@ 4.2 Experiment 1: Predicting Performance among Homogeneous Models
We use 145 models derived from Qwen 2.5-7B, dividing them into 105 models for training, 20 for validation, and 18 for testing. (The detailed configuration of fine-tuned and merged models is listed in the Appendix.) We sample 1,000 instances from the leaderboard benchmarks as the training set, ensuring each model solves all 1,000 instances. This choice follows findings that performance on 100–600 samples reliably predicts overall benchmark performance (Polo et al., 2024; Kipnis et al., 2025).
Table 1: Correlation coefficients between predicted and actual model performance across various benchmarks for a homogeneous set of models derived from Qwen 2.5-7B. Our proposed Lineage-Regularized Matrix Factorization achieves the highest overall correlation (0.719), demonstrating the effectiveness of combining lineage information with collaborative filtering. Note the benchmark-specific patterns: LRMF excels on instruction-following (IFEval) and general reasoning tasks, while Model Lineage Averaging performs better on specialized knowledge (GPQA) and complex reasoning scenarios (MuSR) where performance characteristics tend to be more directly inherited from parent models.
Method all bbh gpqa ifeval math mmlu-pro musr
Lineage-Regularized MF 0.719 0.563 -0.183 0.702 0.324 0.667 0.208
Model Lineage Averaging 0.367 0.412 0.433 0.175 0.175 0.472 0.612
NCF with factors 0.460 0.493 0.310 0.230 0.253 0.506 0.612
@@ Results
Table 1 shows the correlation coefficients between the actual and predicted performance of each LLM across different benchmarks. For overall performance estimation across the entire leaderboard, Lineage-Regularized Matrix Factorization (LRMF) achieved the highest correlation (0.719), followed by NCF with factors (0.460), and Model Lineage Averaging (0.367). These results demonstrate that considering both lineage information and collaborative filtering patterns yields the most accurate predictions.
However, benchmark-specific results reveal notable differences that highlight the distinct characteristics of each method. Model Lineage Averaging performs particularly well on specialized knowledge tasks like GPQA (0.433) and complex reasoning tasks like MuSR (0.612) where it even outperforms the other methods. This suggests that for these particularly challenging benchmarks, fine-tuning or merging tends to produce relatively modest performance changes, making direct ancestor performance a reliable predictor. In contrast, for instruction-following capabilities measured by IFEval, Model Lineage Averaging performs poorly (0.175) while Lineage-Regularized Matrix Factorization excels (0.702). This substantial difference indicates that instruction-following abilities change significantly during fine-tuning or merging, causing direct lineage information alone to be insufficient, whereas the collaborative patterns captured by LRMF provide more relevant signals.
NCF with factors shows intermediate performance across most benchmarks. In this homogeneous experimental setting, where parameter count and model family are identical across all models, the only meaningful factor is the model type (fine-tuned, merged, or other). This limited factor diversity explains why NCF with factors performs relatively well on structural reasoning tasks like BBH (0.493) and knowledge-intensive benchmarks like MMLU-Pro (0.506), but struggles with the more dramatic performance shifts seen in IFEval (0.230).
@@
As shown in Table 2, Lineage-Regularized Matrix Factorization again achieved the highest overall correlation (0.579), followed by Model Lineage Averaging (0.523) and NCF with factors (0.256). Contrary to our expectations, NCF with factors performed worse compared to the other methods, despite the increased diversity of models potentially providing more informative factors in this heterogeneous setting. This underperformance likely stems from the factor importance bias observed in prior work (Zhang et al., 2024), where SHAP values revealed that certain strong factors like model family dominated the predictions. When such strong explanatory factors exist, the model tends to output similar predictions for all models within the same family without capturing finer distinctions, which can artificially inflate correlation in homogeneous settings but fails in more diverse model ecosystems.
The benchmark-specific trends remain largely consistent with Experiment 1. Lineage-Regularized Matrix Factorization continues to excel on BBH (0.612), MATH (0.626), and MMLU-Pro (0.617), while Model Lineage Averaging demonstrates stronger performance on GPQA (0.480) and MuSR (0.379). This consistent pattern across different experimental settings reinforces our understanding of how different capabilities transfer through model lineage.
@@
Notably, compared to Experiment 1, both Lineage-Regularized Matrix Factorization and NCF with factors show a relative decline in performance compared to Model Lineage Averaging. For instance, on IFEval, the gap between LRMF and Model Lineage Averaging narrowed substantially (0.370 vs. 0.373). This suggests that in heterogeneous scenarios with greater model diversity, the latent feature space must capture significantly more complex relationships. Direct lineage information becomes relatively more reliable as embedding-based methods struggle to represent the nuanced differences across diverse architectures and training approaches in a unified latent space.
Table 2: Correlation coefficients between predicted and actual model performance across various benchmarks for a heterogeneous set of models comprising all models available on the Hugging Face Open LLM Leaderboard. Our LRMF consistently achieves higher prediction accuracy compared to baseline methods, highlighting its effectiveness in leveraging lineage information even among diverse architectures and training strategies.
Method all bbh gpqa ifeval math mmlu-pro musr
Lineage-Regularized MF 0.579 0.612 -0.081 0.370 0.626 0.617 0.058
Model Lineage Averaging 0.523 0.525 0.480 0.373 0.592 0.544 0.379
NCF with factors 0.256 0.272 0.222 0.078 0.211 0.294 0.131
@@ 5 Conclusion
Our work demonstrates that lineage relationships—the explicit parent-child connections between derived LLMs—significantly improve performance prediction accuracy. Across experiments with 2,934 models and six major benchmarks, our Lineage-Regularized Matrix Factorization (LRMF) framework achieved up to 7–10 percentage point increases in correlation with actual performance by effectively modeling ancestral connections between models.
@@ Appendix A Detailed Model Statistics
We investigated the lineage distribution within this homogeneous group, discovering that the number of lineage connections roughly follows a logarithmic distribution (Figure 1). Furthermore, we visualized the complete lineage network among the Qwen models (Figure 2), revealing that the model with the most lineage connections is Qwen-2.5-7B-Instruct.
Table 3: Composition of the heterogeneous set used in Experiment 1.

### v2

@@ Abstract
Accurately forecasting the performance of Large Language Models (LLMs) before extensive fine-tuning or merging can substantially reduce both computational expense and development time. Although prior approaches like scaling laws account for global factors such as parameter size or training tokens, they often overlook explicit lineage relationships—i.e., which models are derived or merged from which parents. In this work, we propose a novel Lineage-Regularized Matrix Factorization (LRMF) framework that encodes ancestral ties among LLMs via a graph Laplacian regularizer. By leveraging multi-hop parent–child connections, LRMF consistently outperforms conventional matrix factorization and collaborative filtering methods in both instance-level and benchmark-level performance prediction. Our large-scale study includes 2,934 publicly available Hugging Face models and 21,000+ instances across 6 major benchmarks, showing that the introduction of lineage constraints yields up to 0.15–0.30 higher Pearson correlation coefficients with actual performance compared to baseline methods. Moreover, LRMF effectively addresses the cold-start problem, providing accurate estimates for newly derived or merged models even with minimal data. This lineage-guided strategy thus offers a resource-efficient way to inform hyperparameter tuning, data selection, and model combination in modern LLM development.
@@ 1 Introduction
If one could accurately predict the performance of a proposed LLM variant—even before running large-scale optimization—it would profoundly reduce the search space. Such foresight would streamline decisions about (1) the right model size and data volume for pre-training, (2) fine-tuning hyperparameters and data composition (e.g., domain mix), and (3) whether combining certain model checkpoints is worthwhile. Indeed, prior work has sought to provide partial solutions. Classic scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022) predict an LLM’s loss on the training corpus as a function of parameter size and token count. These laws help developers extrapolate the compute needed to reach a desired perplexity or training loss. Subsequent studies extend these ideas to specific benchmarks, e.g., MMLU (Hendrycks et al., 2021a; Wang et al., 2024), BBH (Suzgun et al., 2023), or MATH (Hendrycks et al., 2021b), by fitting regression models or matrix factorization frameworks to empirical performance data (Zhang et al., 2024; Polo et al., 2024; Kipnis et al., 2025).
@@
We conduct a large-scale empirical study encompassing 2,934 publicly available Hugging Face models and 21,000+ instances across 6 major benchmarks. We demonstrate that incorporating lineage constraints—i.e., explicitly modeling which models were fine-tuned or merged from which parents—consistently boosts predictive accuracy. In fact, compared to standard matrix factorization baselines, our method improves the Pearson correlation coefficient with true performance by up to 0.15–0.30 on certain tasks, affirming that lineage is a key factor in understanding and forecasting LLM behaviors.
Moreover, by leveraging explicit lineage relationships (parent–child relations or merges), our method LRMF tackles the cold-start problem, enabling accurate performance estimates for newly derived or merged models without costly full re-benchmarking. Indeed, our experiments reveal that even with as few as 10 labeled instances per model, we can achieve near state-of-the-art Pearson correlation with fully evaluated baselines—a remarkable improvement over standard collaborative filtering or scaling-based approaches. Taken together, our findings suggest that preserving and exploiting the lineage graph is vital for guiding resource-efficient decisions in LLM development.
@@ 3 Method
Our goal is to predict the performance of Large Language Models (LLMs) on specific tasks or benchmarks by leveraging partial observations and model relationships. In particular, we address the cold-start setting, where we estimate the performance of LLMs that have not yet been constructed or evaluated.
@@ 4.1 Experimental Setting
We compare three methods in our experiments:
Baseline: Neural Collaborative Filtering with Factors (NCF with factors)
Among our baselines, this is the only approach that estimates cold-start performance while explicitly leveraging historical results. We focus on cold-start methods that do not assume any target-model evaluations. [^1] Methods such as Polo et al. (2024) and Kipnis et al. (2025) assume the target model has been evaluated on at least a subset of instances, and are therefore inapplicable under our strict cold-start setting. Likewise, scaling-law approaches (e.g., Ruan et al. (2024)) extrapolate from factors like parameter count or pretraining data volume—quantities that remain essentially unchanged under post-training or model merging—making them tend to be less informative for post-training or model merging for the present objective. We use the same network architecture as prior work (Zhang et al., 2024): a two-layer MLP with 128 hidden units per layer. Owing to practical constraints when collecting data from the leaderboard, we restrict model factors to three attributes: architecture type (50 categories; e.g., qwen2forcasualllm, gemma2forcasuallm), model type (three categories: fine-tuned, merged, and other—including base models), and parameter size. For task factors, we include benchmark identity (six majar benchmarks decomposed into 39 sub-benchmarks). All factors are embedded and concatenated with the model and instance embeddings before being fed to the MLP.
Proposed-1: Model-Lineage Averaging (MLA)
We construct lineage relationships by extracting metadata from Hugging Face model cards. For fine-tuned models, we identify parent models via the base model:finetune: tag; for merged models, we use the base model:merge: tag and, when available, the merge config.json file. To ensure consistency across the corpus, we restrict model factors to those reliably present in the model-card JSON—“architecture type,” “model type,” and “parameter size.” We omit factors such as dataset size or batch size because they are not reported consistently; in the homogeneous setting, such omitted factors are likely similar across models and thus have limited impact on our conclusions.
Proposed-2: Lineage-Regularized Matrix Factorization (LRMF)
We construct the instance-similarity graph from semantic embeddings of the prompts. By default, we use Snowflake/snowflake-arctic-embed-l-v2.0 to embed each prompt, compute cosine similarities, and connect each instance to its top-$k$ nearest neighbors (k=20k{=}20 unless otherwise noted). The model-lineage graph reuses the lineage information described in the Model Lineage Averaging baseline, with an edge indicating a direct derivation between models. The model is implemented in PyTorch and optimized with Adam (learning rate $3\times 10^{-3}$). We train for at most 10,000 epochs with early stopping based on the validation set (patience of 100 epochs).
@@
Hyperparameters $\lambda_{L2}$, $\lambda_{\mathcal{M}}$, and $\lambda_{\mathcal{X}}$ were selected via grid search to maximize AUC-ROC on the development set, yielding the optimal values $\lambda_{L2}=10^{-5}$, $\lambda_{\mathcal{M}}=10^{-4}$, and $\lambda_{\mathcal{X}}=10^{-5}$ (see Appendix B for details).
@@
We evaluate each method using two complementary metrics that assess different aspects of prediction ability:
  • AUC-ROC (Individual Performance Estimation): Measures the ability to accurately estimate whether a single LLM’s performance on a specific task exceeds a threshold, distinguishing correct (1) vs. incorrect (0) predictions at the instance level.
  • Pearson Correlation (Relative Performance Ranking): Measures the ability to correctly preserve the relative performance hierarchy among multiple candidate LLMs on a given task, computed as the Pearson correlation coefficient between predicted and true average scores across benchmarks.
Pearson correlation is crucial for our goal of identifying the most promising model variants without exhaustive benchmarking, as it quantifies how well our method maintains the correct ranking of LLM performance. We report both metrics for each benchmark to provide a comprehensive assessment of individual score estimation accuracy and ranking fidelity.
@@ 4.2 Experiment 1: Predicting Performance among Homogeneous Models
We use 145 models derived from Qwen 2.5-7B, dividing them into 105 models for training, 20 for validation, and 18 for testing. (The detailed configuration of fine-tuned and merged models is listed in the Appendix.) We sample 1,000 instances from the leaderboard benchmarks and ensure that every model is both trained and evaluated on the same set of 1,000 instances. This choice follows findings that performance on 100–600 samples reliably predicts overall benchmark performance (Polo et al., 2024; Kipnis et al., 2025). For this experiment, we repeat the sampling procedure five times with different random seeds and report the mean and standard deviation across these runs.
(a) AUC-ROC
(b) Pearson correlation
Figure 1: Performance evaluation results for a homogeneous set of models derived from Qwen 2.5-7B across various benchmarks. (a) AUC-ROC scores show that all methods achieve similar absolute performance estimation capability (0.83–0.86). (b) Pearson correlation reveals significant differences in relative performance ranking ability, with our proposed Lineage-Regularized MF achieving the highest overall correlation (0.642), substantially outperforming NCF with factors (0.490) and Model Lineage Averaging (0.301). This demonstrates the effectiveness of combining lineage information with collaborative filtering for preserving model performance hierarchies. Note the benchmark-specific patterns: LRMF excels on instruction-following (IFEval) and general reasoning tasks, while Model Lineage Averaging performs better on specialized knowledge (GPQA) where performance characteristics tend to be more directly inherited from parent models.
@@ Results
Figure 1 reports both AUC-ROC and Pearson correlation between the actual and predicted performance of each LLM across different benchmarks. All methods achieve high AUC-ROC scores (approximately 0.83–0.86), with our proposed Lineage-Regularized MF leading marginally. However, when evaluating relative performance via Pearson correlation, LRMF attains a substantially higher score (0.642) compared to NCF with factors (0.490) and Model Lineage Averaging (0.301). This discrepancy indicates that while existing methods can reasonably estimate absolute performance (as reflected by AUC-ROC), they struggle to distinguish the fine-grained performance ordering among multiple candidate models. In contrast, LRMF’s incorporation of lineage information enables accurate prediction of model hierarchies beyond individual performance estimation.
Benchmark-specific results reveal meaningful contrasts in method effectiveness. On IFEval, which measures instruction-following capability, fine-tuning typically induces substantial behavioral changes. Consequently, methods relying primarily on ancestor scores—such as Model Lineage Averaging (MLA)—systematically underperform (correlation: 0.175), whereas LRMF, which exploits collaborative patterns beyond direct lineage, achieves substantially higher accuracy (correlation: 0.702). Conversely, on specialized knowledge benchmarks (GPQA) and complex reasoning tasks (MATH, MuSR), the Qwen-2.5-7B family demonstrates limited performance—most models solve very few problems—making relative ranking inherently challenging and resulting in lower correlations across all methods.
@@
Figure 2 reports both AUC-ROC and Pearson correlation in the heterogeneous setting. All methods again exhibit similarly high AUC-ROC ($\approx$ 0.84–0.88), while Pearson correlation shows clear separation: LRMF reaches 0.615, Model Lineage Averaging 0.504, and NCF with factors 0.301. This mirrors the homogeneous results—while traditional approaches can estimate absolute accuracy, lineage-aware modeling in LRMF most reliably captures the relative ordering of diverse models.
Compared with Experiment 1, NCF with factors shows a relative decline. This underperformance likely reflects the factor-importance bias observed in prior work (Zhang et al., 2024), where SHAP analyses indicated that coarse factors (e.g., model family) dominate predictions. When such strong explanatory factors exist, the model tends to collapse predictions within each family and fails to resolve within-family distinctions—behavior that may inflate correlation in homogeneous settings but does not transfer to heterogeneous ecosystems.
@@
Benchmark-specific trends are largely consistent with Experiment 1. Lineage-Regularized Matrix Factorization continues to excel on BBH, IFEval, MATH, and MMLU-Pro, whereas Model Lineage Averaging remains stronger on GPQA and MuSR. This stable pattern across settings reinforces our view of how distinct capabilities propagate along model lineage. Notably, MLA becomes relatively stronger in this heterogeneous scenario, and the gap between LRMF and MLA narrows. This suggests that greater model diversity requires the latent space to encode more complex relationships; in this regime, direct lineage information serves as a comparatively stable signal, whereas embedding-based methods can struggle to represent nuanced differences across architectures and training recipes within a unified latent space.
(a) AUC-ROC (absolute evaluation)
(b) Pearson correlation (relative evaluation)
Figure 2: Performance evaluation results for heterogeneous models from the Hugging Face Open LLM Leaderboard. (a) AUC-ROC scores remain similar across methods. (b) Pearson correlation shows LRMF achieving the highest overall correlation (0.615), demonstrating effectiveness across diverse architectures and training strategies.
4.4 Experiment 3: Routing
To evaluate the practical applicability of our performance predictions, we consider an instance-level routing scenario where each input must be dynamically assigned to the LLM expected to perform best. We compare five routing strategies: (i) LRMF-based Routing (ours), (ii) Model Lineage Averaging (MLA)-based Routing, (iii) Neural Collaborative Filtering (NCF) with factors-based Routing, (iv) Random Routing, and (v) Best Model, which always assigns all instances to the single model with the highest average true performance.
As shown in Figure 3, LRMF-based routing consistently outperforms all baselines in both homogeneous and heterogeneous settings. Crucially, our method exceeds the Best Model baseline—the minimum threshold for effective instance-level routing—demonstrating that dynamic model assignment based on predicted performance can improve overall benchmark results beyond using a single optimal model.
These findings indicate that explicit modeling of lineage relationships substantially improves model selection in practical routing scenarios, enabling more effective dynamic assignment of inputs to maximize overall performance.
(a) Routing performance in homogeneous setting
(b) Routing performance in heterogeneous setting
Figure 3: Instance-level routing performance across benchmarks in (a) homogeneous and (b) heterogeneous settings. Each test instance is dynamically assigned to the model with the highest predicted score according to different methods. We compare LRMF-based routing against MLA, NCF with factors, Random assignment, and Best Model (always using the single top-performing model per benchmark). LRMF routing consistently outperforms all baselines in both settings, demonstrating that incorporating lineage information enables more effective dynamic model selection that exceeds the performance of using a single optimal model for all instances.
@@ 5 Conclusion
Our work demonstrates that lineage relationships—the explicit parent-child connections between derived LLMs—significantly improve performance prediction accuracy. Across experiments with 2,934 models and six major benchmarks, our Lineage-Regularized Matrix Factorization (LRMF) framework achieved up to 0.15–0.30 increases in correlation with actual performance by effectively modeling ancestral connections between models.
@@ Appendix A Detailed Model Statistics
We investigated the lineage distribution within this homogeneous group, discovering that the number of lineage connections roughly follows a logarithmic distribution (Figure 4(a)). Furthermore, we visualized the complete lineage network among the Qwen models (Figure 5), revealing that the model with the most lineage connections is Qwen-2.5-7B-Instruct.
A.2 Experiment 2: Heterogeneous Models (All Models)
Lineage Analysis of Heterogeneous Models
We also conducted a lineage analysis across the full heterogeneous set, uncovering again a roughly logarithmic distribution in the number of lineage connections (Figure 4(b)). Approximately 60% of these models are either fine-tuned or merged variants, suggesting lineage information is widely applicable for modeling relationships and improving predictions across a broad spectrum of publicly available models.
Table 1: Composition of the homogeneous model set (Qwen 2.5-7B family) used in Experiment 1.